OpenStudy (anonymous):

Find the inclination θ of a line passing through the points (2, -2), (7, 9 Possible Answers A. 1.144 radian(s) or 65.556 degrees B. -1.44 radian(s) or -65.556 degrees C. -1.997 radian(s) or -114.444 degrees D. -0.427 radian(s) or -24.444 degrees E. 0.427 radian(s) or 24.444 degrees

OpenStudy (anonymous):

Exact Duplicate: http://openstudy.com/users/doublebass7#/updates/4fc15a8ce4b0964abc833ee4

OpenStudy (phoenixfire):

you have adjacent and opposite.. so you use tan(x) = o/a a = 7-2 o=9-(-2) x=taninverse(o/a)